OpenStudy (anonymous):

What affect does k = -2 have on the function f(x) = log(x)?

OpenStudy (anonymous):

The graph moves 2 units to the right. The graph moves 2 units to down. The graph moves 2 units to the left. The graph is stretched by a factor of |2|.

OpenStudy (anonymous):

What is k here??

OpenStudy (ash2326):

@coolaidd I think you missed K somewhere, could you please check the question again?

OpenStudy (hba):

this works just like the quadratic form: y = a(x - h)² + k i.e. f(x) = alog(x - h) + k so h moves the curve left and right k moves the curve up and down

OpenStudy (hba):

now i suppose u can tell where does it move

OpenStudy (perl):

man that pic is cheesy